OpenStudy (anonymous):

Find the missing side lenght of the right triangle. Round to the nearest tenth. a = ?,b = 4,c = 41

OpenStudy (anonymous):

sqrt(41^2-4^2)

OpenStudy (anonymous):

=40.8

OpenStudy (dumbcow):

a^2 + b^2 = c^2 a^2 + 4^2 = 41^2 a^2 = 41^2 - 4^2 a = sqrt(41^2 - 4^2) a = 40.8

OpenStudy (anonymous):

thanks, thanks, thanks!!!

OpenStudy (anonymous):

maybe I can sleep now :)

OpenStudy (dumbcow):

haha your welcome

OpenStudy (anonymous):

:D Thanks again!

OpenStudy (anonymous):

what is the time now?

OpenStudy (anonymous):

2:40am

OpenStudy (anonymous):

oh...ok

OpenStudy (anonymous):

good night....math dreams

OpenStudy (anonymous):

what about you?

OpenStudy (anonymous):

i'm not from usa

OpenStudy (anonymous):

so where you from?

OpenStudy (anonymous):

is your time 11:42am?

OpenStudy (anonymous):

India...

OpenStudy (anonymous):

Oh pretty close!

OpenStudy (anonymous):

12:12

OpenStudy (anonymous):

ok...wow

OpenStudy (anonymous):

do u study in school...

OpenStudy (anonymous):

Thank you, I'm trying.

OpenStudy (anonymous):

I'm homeschooled

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

Do you study?

OpenStudy (anonymous):

ok u should go now...

OpenStudy (anonymous):

yeah...obviously

OpenStudy (anonymous):

Ok! Thank you! see you maybe again?! ;)

OpenStudy (anonymous):

yea... nice to talking u

OpenStudy (anonymous):

great, keep it up!!!

OpenStudy (anonymous):

same here. Bye!

OpenStudy (anonymous):

Oh by the way! ap ka naam ka matlab kia hai ? :D

OpenStudy (anonymous):

apka nam ka matlab kya hai...

OpenStudy (anonymous):

great surprise....ha...

OpenStudy (anonymous):

teek ha

OpenStudy (anonymous):

ok take care. Bye

OpenStudy (anonymous):

bye.